concn. of H2SO4 98% by mass, density of soln is 1.80 g/ml calculate vol. of acid required to make 1 lit. of 0.1M H2SO4 soln.

Grade:12

1 Answers

Priyansh Goel
14 years ago

First find molarity of H2SO4 using formula 10xd/M0 where x = mass%, d = density, M0 = molecular mass.

So, Molarity(M) = 10*98*1.8/98=  18.

Now apply molarity equation,M1V1  =  M2V2

18*V = 0.1*1

v = 0.1/18 = 0.005lt